# Network Alignment with Transferable Graph Autoencoders

### v2

@@ 5.3.2 Transferability across perturbation models
[…]
Hit@1
Hit@5
Hit@10
Hit@50
Hit@1
Hit@5
Hit@10
Hit@50
Netsimile
2.59%
8.32%
12.09%
26.42%
1.07%
2.77%
4.74%
15.03%
Spectral
1.40%
4.62%
7.21%
16.34%
0.54%
1.34%
2.95%
13.95%
GAE
0.83%
3.28%
5.43%
10.51%
0.27%
0.81%
1.97%
8.94%
$\text{WAlign}_{\text{final}}$
29.21%
50.28%
59.01%
74.05%
22.91%
46.60%
52.42%
46.51%
$\text{WAlign}_{\text{best}}$
62.02%
81.96%
87.31%
93.89%
$\mathbf{36.40\%}$
$\mathbf{53.94\%}$
$\mathbf{67.08\%}$
85.33%
[…]
73.18%
90.97%
94.70%
98.04%
23.79%
45.08%
56.17%
83.18%
[…]
$\mathbf{73.48\%}$
$\mathbf{91.73\%}$
$\mathbf{94.79\%}$
$\mathbf{98.17\%}$
34.18%
49.37%
56.89%
$\mathbf{85.36\%}$
Figure 4: Graph Matching accuracy of T-GAE when trained on different perturbation models
In the previous experiments, we showed the transferability of T-GAE across different datasets. In this set of experiments, we test the robustness of the proposed method by training and testing on different perturbation models. We consider two perturbation models:
We consider two scenarios: i) We train the T-GAE encoder on augmented samples generated from uniform perturbation model, where we add or remove edges randomly, and we test on different graphs that are generated from the same perturbation model (uniform). ii) We train the T-GAE encoder on augmented samples generated from the degree perturbation model, where the edges incident to nodes that have higher degree are more probable to be edited, and we test on graphs generated from a different perturbation model (uniform). The results are illustrated in Figure 4.
We observe that when training and testing is performed with graphs generated from different perturbation models, the performance of T-GAE is accurate and stable. This is concrete evidence for the robustness of T-GAE in matching graphs from distribution-shifted data sources.
5.4 Sub-graph Matching Experiments
@@
[…]
We observe a significant improvement in matching accuracy with GNN-based methods compared to traditional graph or node embedding techniques. These results demonstrate the ability of GNNs to generate expressive and robust node embeddings compared to classical algorithms. In particular, our proposed framework, T-GAE, consistently achieves the best final performance among all competing methods. This suggests that the training framework (8), illustrated in Fig. 2, provides an efficient approach to network alignment. It is also notable, that T-GAE is very stable during training as the final graph matching accuracy converges to the best result it achieved. This result indicates that the proposed training framework is robust to problems like over-fitting, not sensitive to hype-parameters shift and is capable to be extended to different types of networks and graph matching tasks.
@@ Appendix A Notation and further reading
$\mathcal{G}$triangleq
Graph
$\mathcal{V}$triangleq
Set of nodes
$\mathcal{E}$triangleq
Set of edges
$N$triangleq
Number of nodes
$\bm{D}$triangleq
Degree matrix
$\bm{S}$triangleq
$\{0,1\}^{N\times N}$ adjacency matrix
$\bm{X}$triangleq
$N\times D$ feature matrix
$\bm{H}$triangleq
aggregation results of GNN convolution
$\bm{W}$triangleq
weight matrix of the Graph Neural Network
$\bm{N}_{v}$triangleq
neighbors of node v
$\bm{I}$triangleq
Identity matrix
$\bm{0}$triangleq
vector or matrix of zeros
$\bm{A}^{T}$triangleq
transpose of matrix $\bm{A}$
$\bm{A}_{rc}$triangleq
entry at r-th row and j-th column of matrix $\bm{A}$
$\lVert\cdot\rVert_{F}$triangleq
Frobenius norm
Graph matching, as an important problem, has been studied not only by the community of data mining, it has also been mathematically and statistically investigated. A number of approaches have been proposed to solve this problem for Erdős Rényi random graphs $G(n,\frac{d}{n})$. It has been proved that a perfectly true vertex correspondence can be recovered in polynomial time with high probability (Ding et al., 2020). Furthermore, a sharp threshold has been proved for both Erdős Rényi model and Gaussian model (Wu et al., 2022). Most recently, a novel approach to calculate similarity scores based on counting weighted trees rooted at each vertex has been proposed (Mao et al., 2023). Such approach has been proved to be effective in solving the aforementioned network alignment problem on random graphs with high probability. Readers are encouraged to refer to the authors of these publications (Ding et al., 2020; Wu et al., 2022; Mao et al., 2023) for further reading.
@@ Appendix B Spectral characterization of GNNs
where $\lambda_{n},\bm{v}_{n}$ are the $n-$th eigenvalue and eigenvector and $a_{n}^{(i)}=\bm{v}_{n}^{T}\bm{X}^{(l)}\sum_{k=0}^{K-1}\lambda_{n}^{k}\bm{H}_{k}^{(l)}[:,i]$ is a scalar related to the Graph Fourier Transform (GFT) of $\bm{X}^{(l)}$ (Sardellitti et al., 2017). It is clear from equation (11) that the output of each layer is a linear combination of the adjacency eigenvectors, followed by a pointwise non-linearity. Thus, a GNN can produce unique and more powerful graph embeddings than spectral methods by processing the eigenvectors and eigenvalues of the adjacency matrix.
@@ Appendix E Implementation Details
In this section we discuss the implementation details of our framework. The code and datasets used in all experiments can be found here: https://github.com/GraphMatching/Graph-Matching
@@ E.1 Assignment Optimization
The problem in (40) can be optimally solved in $\mathcal{O}\left(N^{3}\right)$ flops by the Hungarian algorithm (Kuhn, 1955a). To avoid this computational burden we employ the greedy Hungarian approach that has computational complexity $\mathcal{O}\left(N^{2}\right)$ and usually works well in practice.
@@
$\bm{P}~{}:=\bm{0}_{N\times N}$   // Initialize permutation matrix
$\bm{D}~{}:=$ PairwiseDistance $\left(\bm{X},\hat{\bm{X}}\right)$   // pairwise Euclidean distance
@@
for $n$=1 to $N$ do
@@ E.2 Datasets
  • Celegans (Kunegis, 2013): The vertices represent proteins and the edges their protein-protein interactions.
  • Arenas Email (Leskovec & Krevl, 2014): The email communication network at the University Rovira i Virgili in Tarragona in the south of Catalonia in Spain. Nodes are users and each edge represents that at least one email was sent.
  • Douban (Zhang & Tong, 2016): Contains user-user relationship on the Chinese movie review platform. Each edge implies that two users are contacts or friends.
  • Cora (Sen et al., 2008): The dataset consists of 2708 scientific publications, with edges representing citation relationships between them. Cora has been one of the major benchmark datasets in many graph mining tasks.
  • Dblp (Pan et al., 2016): A citation network dataset that is extracted from DBLP, Association for Computing Machinery (ACM), Microsoft Academic Graph (MAG), and other sources. It is considered a benchmark in multiple tasks.
  • Coauthor_CS (Shchur et al., 2018): The coauthorship graph is generated from MAG. Nodes are the authors and they are connected with an edge if they coauthored at least one paper.
  • ACMDBLP (Zhang & Tong, 2019): The citation networks that share some common nodes. The task is to identify the publications that appear in both networks.
  • Douban OnlineOffline (Zhang & Tong, 2016): The two social networks contained in this dataset represents the online and offline events of the Douban social network. The task is to identify users that participate in both online and offline events.
@@ E.3.1 Graph Neural Network(GNN) based methods
To have a fair comparision with the node embedding models, all GNN methods are set to have the hidden dimension equal to 4 in the graph matching experiments. In sub graph matching, the hidden dimension of GNNS are fine-tuned from 4 to 512. Other parameters are the same as the author suggested.
  • WAlign (Gao et al., 2021) fits a GNN to each of the input graphs, trains the model by reconstructing the given inputs and minimizing an approximation of Wasserstein distance between the node embeddings. We use the author’s implementation from https://github.com/gaoji7777/walign.git.
  • GAE, VGAE(Kipf & Welling, 2016) are self-supervised graph learning frameworks that are trained by reconstructing the graph. The encoder is a GCN(Kipf & Welling, 2017) and linear decoder is applied to predict the original adjacency. In VGAE, Gausian Noise is introduced to the node embeddings before passing to the decoder. We use the implementation from https://github.com/DaehanKim/vgae_pytorc.
@@ E.3.2 Graph/Node embedding techniques
  • NetSimile (Berlingerio et al., 2013) uses the structural features described earlier to match the nodes of the graphs. Since the NetSimile features are used as input to the T-GAE, they provide a measure to assess the benefit of using T-GAE for node embedding. It proposed 7 egonet-based features, to measure network similarity. We process these features by Algorithm 1 to perform network alignment. The 7-dimensional Netsimile features are:
@@
  • Spectral (Umeyama, 1988) It solves the following optimization problem:
@@
  • DeepWalk (Perozzi et al., 2014): A node embedding approach, simulates random walks on the graph and apply skip-gram on the walks to generate node embedding. We use the implementation from Karateclub https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/neighbourhood/deepwalk.pyThe algorithm is implemented with the default parameters as suggested by this repository, the number of random walks is 10 with each walk of length 80. The dimensionality of embedding is set to be 128. We run the algorithm with 1 epoch and set the learning rate to be 0.05.
  • Node2Vec (Grover & Leskovec, 2016b): An improve version of DeepWalk, it has weights on the randomly generated random walks, to make the neighborhood preserving objective more flexible. We use the implementation from https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/neighbourhood/node2vec.py. The default parameters are used. We simulate 10 random walks on the graph with length 80. p and q are both equal to 1. Dimensionality of embeddings is set to be 4 and we run 1 epoch with learning rate 0.05.
  • GraphWave (Donnat et al., 2018): The structure information of the graphs is captured by simulating heat diffusion process on them. We use the implementation from (Donnat et al., 2018):https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/structural/graphwave.py with the default parameters: number of evaluation points is 200, step size is 0.1, heat coefficient is 1.0 and Chebyshev polynomial order is set to be 100. Note that this implementation does not work on graphs with more than 10,000 nodes, so we exclude this model on the DBLP and Coauthor_CS dataset.
  • LINE (Tang et al., 2015): An optimization based graph embedding approach that aims to preserve local and global structures of the network by considering substructures and structural-quivariant nodes. We use the PyTorch implementation from https://github.com/zxhhh97/ABot. All parameters are set to default as the authors suggested.
@@ E.3.3 Optimization based graph matching algorithms
  • FINAL (Zhang & Tong, 2016) is an optimization approach, following an alignment consistency principle, and tries to match nodes with similar topology. In the case of graphs without attributes FINAL is equivalent to the popular
  • Isorank (Singh et al., 2008) algorithm, whereas using NetSimile as an input to FINAL resulted in inferior performance and was therefore omitted. We use the code in https://github.com/sizhang92/FINAL-KDD16 with $\bm{H}$ being the degree similarity matrix, $\alpha=0.8,~{}\text{maxiter}=30,~{}\text{tol}=1e-4$ as suggested in the repository.
  • ConeAlign (Chen et al., 2020) is a graph embedding based approach. The matching is optimized in each iteration by the Wasserstein Procrustes distances between the matched embeddings calculated on a mini batch in order to preserve scalability. We use the official implementation from https://github.com/GemsLab/CONE-Align and preserved all the suggested parameters.
  • S-GWL (Xu et al., 2019a) matches two given graphs by retrieving node correspondence from the optimal transport associated with the Gromov-Wasserstein discrepancy between the graphs. We use the implementation by the authors in https://github.com/HongtengXu/s-gwl with all the suggested parameters.

### v3

@@ 5.4 Sub-graph Matching Experiments
Hit@1 Hit@5 Hit@10 Hit@50 Hit@1 Hit@5 Hit@10 Hit@50
Netsimile 2.59% 8.32% 12.09% 26.42% 1.07% 2.77% 4.74% 15.03%
Spectral 1.40% 4.62% 7.21% 16.34% 0.54% 1.34% 2.95% 13.95%
GAE 8.1% 22.5% 30.1% 45.1% 3.3% 9.2% 14.1% 32.1%
WAlign 62.02% 81.96% 87.31% 93.89% ${36.40\%}$ ${53.94\%}$ ${67.08\%}$ 85.33%
T-GAE $\mathbf{73.89\%}$ $\mathbf{91.73\%}$ $\mathbf{95.33\%}$ $\mathbf{98.22\%}$ $\mathbf{36.94\%}$ $\mathbf{60.64\%}$ $\mathbf{69.77\%}$ $\mathbf{89.62\%}$
@@
This set of experiments is important in proving the robustness and generalizability of the proposed T-GAE framework for the following reason: (1) As we can observe in Table 3, most optimization methods work well on matching the clean graphs. However, their optimization objective usually prevents them from matching graphs with different numbers of nodes, which is the most realistic scenario in real-world applications. (2) Matching two different graphs is considered challenging for the unknown noise distribution, especially for unsupervised approaches.
We observe a significant improvement in matching accuracy with GNN-based methods (TGAE and WAlign) compared to traditional graph or node embedding techniques. These results demonstrate the ability of GNNs to generate expressive and robust node embeddings compared to classical algorithms. In particular, our proposed framework, T-GAE, consistently achieves the best performance among all competing methods. This suggests that the encoder model illustrated in Fig. 1 and the training framework (8) illustrated in Fig. 2, provide an efficient approach to generate powerful node embedding, that is robust to real-world noise distributions for the task of network alignment.
@@ Appendix A Notation and further reading
$\mathcal{G}$triangleq Graph
$\mathcal{V}$triangleq Set of nodes
$\mathcal{E}$triangleq Set of edges
$N$triangleq Number of nodes
$\bm{D}$triangleq Degree matrix
$\bm{S}$triangleq $\{0,1\}^{N\times N}$ adjacency matrix
$\bm{X}$triangleq $N\times D$ feature matrix
$\bm{H}$triangleq aggregation results of GNN convolution
$\bm{W}$triangleq weight matrix of the Graph Neural Network
$\bm{N}_{v}$triangleq neighbors of node v
$\bm{I}$triangleq Identity matrix
$\bm{0}$triangleq vector or matrix of zeros
$\bm{A}^{T}$triangleq transpose of matrix $\bm{A}$
$\bm{A}_{rc}$triangleq entry at r-th row and j-th column of matrix $\bm{A}$
$\lVert\cdot\rVert_{F}$triangleq Frobenius norm
Graph matching, as an important problem, has been studied not only by the community of data mining, it has also been mathematically and statistically investigated. A number of approaches have been proposed to solve this problem for Erdős Rényi random graphs $G(n,\frac{d}{n})$. It has been proved that a perfectly true vertex correspondence can be recovered in polynomial time with high probability [Ding et al., 2020]. Furthermore, a sharp threshold has been proved for both Erdős Rényi model and Gaussian model [Wu et al., 2022]. Most recently, a novel approach to calculate similarity scores based on counting weighted trees rooted at each vertex has been proposed Mao et al. [2023]. Such approach has been proved to be effective in solving the aforementioned network alignment problem on random graphs with high probability. Readers are encouraged to refer to the authors of these publications [Ding et al., 2020, Wu et al., 2022, Mao et al., 2023] for further reading.
@@ Appendix B Spectral characterization of GNNs
where $\lambda_{n},\bm{v}_{n}$ are the $n-$th eigenvalue and eigenvector and $a_{n}^{(i)}=\bm{v}_{n}^{T}\bm{X}^{(l)}\sum_{k=0}^{K-1}\lambda_{n}^{k}\bm{H}_{k}^{(l)}[:,i]$ is a scalar related to the Graph Fourier Transform (GFT) of $\bm{X}^{(l)}$ [Sardellitti et al., 2017]. It is clear from equation (11) that the output of each layer is a linear combination of the adjacency eigenvectors, followed by a pointwise non-linearity. Thus, a GNN can produce unique and more powerful graph embeddings than spectral methods by processing the eigenvectors and eigenvalues of the adjacency matrix.
@@ Appendix E Implementation Details
In this section we discuss the implementation details of our framework. The code and datasets used in all experiments can be found here: https://github.com/GraphMatching/Network-Alignment-with-Transferable-Graph-Autoencoders
@@ E.1 Assignment Optimization
The problem in (28) can be optimally solved in $\mathcal{O}\left(N^{3}\right)$ flops by the Hungarian algorithm [Kuhn, 1955a]. To avoid this computational burden we employ the greedy Hungarian approach that has computational complexity $\mathcal{O}\left(N^{2}\right)$ and usually works well in practice.
@@
$\bm{P}~{}:=\bm{0}_{N\times N}$
  // Initialize permutation matrix
$\bm{D}~{}:=$ PairwiseDistance $\left(\bm{X},\hat{\bm{X}}\right)$
  // pairwise Euclidean distance
@@
1 for $n$=1 to $N$ do
2       $i,~{}j:=\text{argmin}~{}(\bm{D})$
3       $r~{}:=\text{rows}~{}[i]$
4       $c~{}:=\text{cols}~{}[j]$
5       $\bm{P}_{rc}:=1$
6       Remove $r$ from rows
7       Remove $c$ from cols
8       Remove the $i$-th row from $\bm{D}$
9       Remove the $j$-th column from $\bm{D}$
10
11return $\bm{P}$
@@ E.2 Datasets
  • Celegans [Kunegis, 2013]: The vertices represent proteins and the edges their protein-protein interactions.
  • Arenas Email [Leskovec and Krevl, 2014]: The email communication network at the University Rovira i Virgili in Tarragona in the south of Catalonia in Spain. Nodes are users and each edge represents that at least one email was sent.
  • Douban [Zhang and Tong, 2016]: Contains user-user relationship on the Chinese movie review platform. Each edge implies that two users are contacts or friends.
  • Cora [Sen et al., 2008]: The dataset consists of 2708 scientific publications, with edges representing citation relationships between them. Cora has been one of the major benchmark datasets in many graph mining tasks.
  • Dblp [Pan et al., 2016]: A citation network dataset that is extracted from DBLP, Association for Computing Machinery (ACM), Microsoft Academic Graph (MAG), and other sources. It is considered a benchmark in multiple tasks.
  • Coauthor_CS [Shchur et al., 2018]: The coauthorship graph is generated from MAG. Nodes are the authors and they are connected with an edge if they coauthored at least one paper.
  • ACM--DBLP [Zhang and Tong, 2019]: The citation networks that share some common nodes. The task is to identify the publications that appear in both networks.
  • Douban Online--Offline [Zhang and Tong, 2016]: The two social networks contained in this dataset represents the online and offline events of the Douban social network. The task is to identify users that participate in both online and offline events.
@@ E.3.1 Graph Neural Network(GNN) based methods
To have a fair comparison with the node embedding models, GAE is implemented using the same set of parameters as T-GAE, which can be found at Section E.4. WAlign is implemented using parameters suggested by the author. We report the best results they achieved during training.
  • WAlign [Gao et al., 2021] fits a GNN to each of the input graphs, trains the model by reconstructing the given inputs and minimizing an approximation of Wasserstein distance between the node embeddings. We use the author’s implementation from https://github.com/gaoji7777/walign.git.
  • GAE, VGAE[Kipf and Welling, 2016] are self-supervised graph learning frameworks that are trained by reconstructing the graph. The encoder is a GCN[Kipf and Welling, 2017] and linear decoder is applied to predict the original adjacency. In VGAE, Gausian Noise is introduced to the node embeddings before passing to the decoder. We use the implementation from https://github.com/DaehanKim/vgae_pytorch. We train GAE by reconstructing the given network using the netsimile node embedding.
@@ E.3.2 Graph/Node embedding techniques
  • NetSimile [Berlingerio et al., 2013] uses the structural features described earlier to match the nodes of the graphs. Since the NetSimile features are used as input to the T-GAE, they provide a measure to assess the benefit of using T-GAE for node embedding. It proposed 7 egonet-based features, to measure network similarity. We process these features by Algorithm 1 to perform network alignment. The 7-dimensional Netsimile features are:
@@
  • Spectral [Umeyama, 1988] It solves the following optimization problem:
@@
  • DeepWalk [Perozzi et al., 2014]: A node embedding approach, simulates random walks on the graph and apply skip-gram on the walks to generate node embedding. We use the implementation from Karateclub https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/neighbourhood/deepwalk.pyThe algorithm is implemented with the default parameters as suggested by this repository, the number of random walks is 10 with each walk of length 80. The dimensionality of embedding is set to be 128. We run the algorithm with 1 epoch and set the learning rate to be 0.05.
  • Node2Vec [Grover and Leskovec, 2016b]: An improve version of DeepWalk, it has weights on the randomly generated random walks, to make the neighborhood preserving objective more flexible. We use the implementation from https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/neighbourhood/node2vec.py. The default parameters are used. We simulate 10 random walks on the graph with length 80. p and q are both equal to 1. Dimensionality of embeddings is set to be 4 and we run 1 epoch with learning rate 0.05.
  • GraphWave [Donnat et al., 2018]: The structure information of the graphs is captured by simulating heat diffusion process on them. We use the implementation from [Donnat et al., 2018]:https://github.com/benedekrozemberczki/karateclub/blob/master/karateclub/node_embedding/structural/graphwave.py with the default parameters: number of evaluation points is 200, step size is 0.1, heat coefficient is 1.0 and Chebyshev polynomial order is set to be 100. Note that this implementation does not work on graphs with more than 10,000 nodes, so we exclude this model on the DBLP and Coauthor_CS dataset.
  • LINE [Tang et al., 2015]: An optimization based graph embedding approach that aims to preserve local and global structures of the network by considering substructures and structural-quivariant nodes. We use the PyTorch implementation from https://github.com/zxhhh97/ABot. All parameters are set to default as the authors suggested.
@@ E.3.3 Optimization based graph matching algorithms
  • FINAL [Zhang and Tong, 2016] is an optimization approach, following an alignment consistency principle, and tries to match nodes with similar topology. In the case of graphs without attributes FINAL is equivalent to the popular
  • Isorank [Singh et al., 2008] algorithm, whereas using NetSimile as an input to FINAL resulted in inferior performance and was therefore omitted. We use the code in https://github.com/sizhang92/FINAL-KDD16 with $\bm{H}$ being the degree similarity matrix, $\alpha=0.8,~{}\text{maxiter}=30,~{}\text{tol}=1e-4$ as suggested in the repository.
  • ConeAlign [Chen et al., 2020] is a graph embedding based approach. The matching is optimized in each iteration by the Wasserstein Procrustes distances between the matched embeddings calculated on a mini batch in order to preserve scalability. We use the official implementation from https://github.com/GemsLab/CONE-Align and preserved all the suggested parameters.
  • S-GWL [Xu et al., 2019a] matches two given graphs by retrieving node correspondence from the optimal transport associated with the Gromov-Wasserstein discrepancy between the graphs. We use the implementation by the authors in https://github.com/HongtengXu/s-gwl, since the performance of S-GWL is very sensitive to the parameter gamma, as suggested by the authors, we fine-tuned this parameter over the range of $[0.001,0.1]$ for each dataset on the cleaned graph, and use that optimal parameter for all other experiments on this dataset.
@@ E.4 T-GAE Parameters
We implement our proposed Transferable graph autoencoder (T-GAE) framework, following Equations (6), (7), (8), for the tasks described later in the section.
For the graph matching experiments in Tables 3 and 5. We use 1 MLP and 8 GNN layers, followed by another MLP to design the Encoder $\phi\left(\bm{X},{\bm{S}};\mathcal{H}\right)$ of T-GAE, as illustrated in Fig. 1. For the GNN layers we use graph isomorphism network (GIN) [Xu et al., 2019b], which is described by (4) with $f$ being the summation function, and $g$ being the MLP. For both graph matching and sub-graph matching experiments, we use hidden dimension of 16 and output dimension of 8. To ensure the fairness of comparison, this set of parameters are tuned on matching the cleaned version of Celegans, no other fine-tuning is conduced for each individual dataset or perturbation levels.
For the sub-graph matching experiments in Table 4, on ACMDBLP, we use 1 MLP and 12 GNN layers, followed by another MLP as encoder, the hidden and output dimension is set to 1024. On Douban OnlineOffline, we use 6 GNN layers with hidden and output dimensions equal to 512. This set of parameters is chosen based on the best matching accuracy achieved.
To summarize, detailed settings of our experiments are:
  • Optimizer, epoch and learning rate: We train our model using Adam optimizer for 200 epochs and initial learning rate equal to $0.001$ for all graph matching experiments. For sub-graph matching experiments, we train the model using Adam optimizer for 50 epochs and the initial learning rate is set to $0.0001$.
  • Dimensions: The hidden dimension of the GNN modules is equal to $16$ and that of the output layer (encoder) is equal to 8 for all graph matching experiments. For sub-graph matching experiments, on ACMDBLP, the hidden and output dimension equal to 1024. On Douban OnlineOffline, they are both set to 512.
  • Number of layers: On all graph matching experiments, the number of GNN layers equals to 8, and for sub-graph matching experiments, we use 12 GNN layers for ACMDBLP and 6 GNN layers for Douban OnlineOffline.
  • Activation function: The activation function for each GNN layer is ReLU.
  • Decoder: We use an 1-layer decoder followed by the sigmoid function.
  • Data Augmentation: To conduct the experiments in Tables 3 and 5, we augment the dataset with $10$ randomly generated perturbed graphs for each perturbation level and perturbation model.
  • Compute Resources: All our experiments are conduced on AMD Ryzen Threadripper 3960X 24-Core Processor and Nvidia RTX A6000 GPU.
Appendix F Efficiency and Run-time Comparison
We analyze the efficiency of the proposed graph matching framework by comparing its running time with the competing algorithms. The GNN-based methods achieve at most $\times$50 less running time on both graph matching and subgraph matching tasks. The proposed training objective (8) scales well from networks with 1,000 nodes [Zhang and Tong, 2016] to denser networks with $\times$10 nodes, with minor running time increase, compared to other node embedding techniques.
We also noticed that compared to the other GNN based framework WAlign Gao et al. [2021], T-GAE takes more time to train on the larger scale networks as illustrated in Figure 5. This is because the deep GNN is trained by the structure of the network. The training takes more time as the netowrk scales up and the number of nodes and edges grow dramatically, whereas the training labels for WAlign come from the input feature, which does not scale up with as the size of the training graphs increase. This could be one of the limitations to the proposed training paradigm and that is why we leverage transfer learning to efficiently match large scale graphs, as illustrated in Table3.
Algorithm Celegans Arenas Douban Cora Dblp Coauthor CS
T-GAE 0.0980.098{0.098}0.098 0.2410.241{0.241}0.241 1.9341.934{1.934}1.934 0.8420.842{0.842}0.842 173.836173.836{173.836}173.836 193.289193.289{193.289}193.289
GAE 0.0940.094{0.094}0.094 0.2520.252{0.252}0.252 1.9491.949{1.949}1.949 0.8570.857{0.857}0.857 174.410174.410{174.410}174.410 194.062194.062{194.062}194.062
WAlign 0.0780.078{0.078}0.078 0.2050.205{0.205}0.205 1.8001.800{1.800}1.800 0.7660.766{0.766}0.766 169.694169.694{169.694}169.694 189.032189.032{189.032}189.032
Spectrul 5.7125.712{5.712}5.712 2.8192.819{2.819}2.819 60.29860.298{60.298}60.298 54.77054.770{54.770}54.770 > 48 hours > 48 hours
GraphWave 8.3088.308{8.308}8.308 32.99432.994{32.994}32.994 281.629281.629{281.629}281.629 131.230131.230{131.230}131.230 5470.7245470.724{5470.724}5470.724 6291.3686291.368{6291.368}6291.368
ConeAlign 1.3331.333{1.333}1.333 3.5003.500{3.500}3.500 31.95531.955{31.955}31.955 13.79913.799{13.799}13.799 887.090887.090{887.090}887.090 1099.1451099.145{1099.145}1099.145
S-GWL 27.84427.844{27.844}27.844 37.44337.443{37.443}37.443 3394.5223394.522{3394.522}3394.522 311.201311.201{311.201}311.201 > 48 hours > 48 hours
FINAL 0.0300.030{0.030}0.030 0.0810.081{0.081}0.081 1.0071.007{1.007}1.007 0.4980.498{0.498}0.498 86.78886.788{86.788}86.788 118.065118.065{118.065}118.065
Appendix G More Baseline Results
Table 3: Graph matching accuracy on 10 randomly perturbed samples under different levels of edge editing for VGAE, LINE and DeepWalk.
Dataset VGAE LINE DeepWalk
$0\%$ Celegans ${0.3\pm 0.1}$ ${1.0\pm 0.5}$ ${1.8\pm 0.6}$
Arenas ${0.1\pm 0.1}$ ${0.2\pm 0.1}$ ${0.3\pm 0.2}$
Douban ${0.0\pm 0.0}$ ${0.0\pm 0.0}$ ${0.1\pm 0.0}$
Cora ${0.1\pm 0.0}$ ${0.0\pm 0.0}$ ${0.1\pm 0.0}$
$1\%$ Celegans ${0.3\pm 0.1}$ ${1.0\pm 0.4}$ ${1.2\pm 0.5}$
Arenas ${0.1\pm 0.1}$ ${0.1\pm 0.1}$ ${0.3\pm 0.1}$
Douban ${0.0\pm 0.0}$ ${0.0\pm 0.0}$ ${0.1\pm 0.0}$
Cora ${0.1\pm 0.1}$ ${0.1\pm 0.0}$ ${0.2\pm 0.1}$
$5\%$ Celegans ${0.6\pm 0.3}$ ${0.9\pm 0.3}$ ${1.0\pm 0.3}$
Arenas ${0.2\pm 0.1}$ ${0.2\pm 0.2}$ ${0.2\pm 0.1}$
Douban ${0.0\pm 0.0}$ ${0.0\pm 0.0}$ ${0.0\pm 0.0}$
Cora ${0.1\pm 0.0}$ ${0.1\pm 0.0}$ ${0.1\pm 0.0}$
Hit rate GraphWave
Hit@1 0.090.09{0.09}0.09
Hit@5 0.360.36{0.36}0.36
Hit@10 0.810.81{0.81}0.81
Hit@50 4.744.74{4.74}4.74
Hit@100 9.129.12{9.12}9.12
Appendix H Degree Perturbation Model Results
Table 5: Graph matching accuracy on 10 randomly perturbed samples under different levels of edge removal on Degree model. The proposed T-GAE is trained on the clean Celegans, Arena, Douban, and Cora networks, and tested on noisy versions of them and the larger Dblp, and Coauthor CS. Accuracy above 80% is highlighted in green, 60% to 80% accuracy is in yellow, and performance below 60% is in red.